# Efficient 3D-Aware Facial Image Editing via Attribute-Specific Prompt Learning

### v1

@@ Abstract
Drawing upon StyleGAN’s expressivity and disentangled latent space, existing 2D approaches employ textual prompting to edit facial images with different attributes. In contrast, 3D-aware approaches that generate faces at different target poses require attribute-specific classifiers, learning separate model weights for each attribute, and are not scalable for novel attributes. In this work, we propose an efficient, plug-and-play, 3D-aware face editing framework based on attribute-specific prompt learning, enabling the generation of facial images with controllable attributes across various target poses. To this end, we introduce a text-driven learnable style token-based latent attribute editor (LAE). The LAE harnesses a pre-trained vision-language model to find text-guided attribute-specific editing direction in the latent space of any pre-trained 3D-aware GAN. It utilizes learnable style tokens and style mappers to learn and transform this editing direction to 3D latent space. To train LAE with multiple attributes, we use directional contrastive loss and style token loss. Furthermore, to ensure view consistency and identity preservation across different poses and attributes, we employ several 3D-aware identity and pose preservation losses. Our experiments show that our proposed framework generates high-quality images with 3D awareness and view consistency while maintaining attribute-specific features. We demonstrate the effectiveness of our method on different facial attributes, including hair color and style, expression, and others. Code : https://github.com/VIROBO-15/Efficient-3D-Aware-Facial-Image-Editing.
@@ 1 Introduction
**footnotetext: Authors contributed equally. Contact information for Amandeep is available on his page, and for Awais, on his page.
StyleGAN [21, 22] has demonstrated exceptional capabilities in generating unconditional, photorealistic 2D images. The disentangled properties present in the learned latent space of StyleGAN have facilitated its utilization for attribute-specific realistic image  [1, 36, 15, 41]. These models edit specified attributes of an input image by finding and modifying semantic editing directions in the latent space of StyleGANs. However, finding an editing direction is tedious and non-intuitive as it requires training classifiers. To circumvent these issues, vision-language foundational models (e.g., CLIP  [32]) have also been used to facilitate text-guided attribute editing [30].
@@
Building on 2D GANs, some works have introduced attribute editing methods to manipulate a set of attributes in 3D [27, 39, 34, 24, 50, 7, 14, 27, 39]. For instance, GMPI [50] modifies 2D StyleGAN by introducing an alpha branch to learn alpha maps and efficiently generate 3D-aware multiplane images. StyleNeRF [14] integrates the neural radiance field (NeRF) into a style-based generator and EG3D [7] which incorporate triplane representation along with style mapping network to generate view consistent images.
However, 3D manipulation methods are constrained to editing a predetermined set of attributes, as they rely on pre-trained attribute classifiers for each new attribute. In contrast, textual prompts offer a more natural and flexible approach, significantly extending the range of manipulations and providing a more intuitive interaction that closely resembles human communication [30]. Nonetheless, compared with 2D, achieving 3D-aware text-guided image editing presents a unique set of challenges. These include translating prompt instructions into semantic editing directions for the generative model’s latent space while maintaining consistency between the edited and original images, ensuring view consistency across different poses, and preserving facial identity.
@@
In this work, we introduce a plug-and-play module designed to efficiently integrate text-guided editing of novel attributes into 3D-aware GANs while preserving both the 3D pose and identity. Our proposed module, Learned Attribute Editor (LAE), finds semantic editing directions specified by a natural language prompt, facilitating text-guided manipulation of novel attributes. The language prompt specifying the target attribute is concatenated with learnable style tokens and passed through a frozen CLIP text encoder. These prompt encodings are passed through a set of style mappers that learn to transform text-specified modifications to the latent space of 3D-aware GAN. The combination of learnable style tokens and vision-language models makes the editing process extremely efficient (from hours to a few minutes), as LAE does not need to train classifiers on large datasets for every new attribute.
@@
Training the LAE module with language supervision poses challenges, as standard CLIP loss [32] may struggle to converge in the 3D latent space. To overcome this issue, we introduce a 3D-aware DCLIP loss [12] which utilizes multi-view images generated on-the-fly. To improve the view consistency and preserve multi-view identity, we introduce novel losses, including style token contrastive loss, IDVC, and several 3D-aware pose and identity preservation losses. These losses enable maintaining the distinctive characteristics of each attribute, aligning the CLIP-space direction with learnable text features and the multi-view generated images, and ensuring consistency across different camera poses. We primarily employ GMPI [50], which adapts StyleGAN to generate color images and corresponding alpha maps. We also demonstrate our method’s plug-and-play nature by integrating it with several 3D-aware GANs. Our experimental results validate the efficacy of our proposed method, showcasing its superior generative performance in both qualitative and quantitative ways. We present a representative sample of editing by our method at random camera poses in Figure 1.
@@ 2 Related Works
Generative 3D-aware image synthesis and manipulation. Extending the capabilities of 2D GANs to 3D settings has gained significant attention recently. These methods rely on a combination of 3D-structure aware inductive bias in the generator and utilization of neural rendering engines to get view-consistent results. These approaches include Mesh-based appraoches [26, 40], Voxel-based GANs [11, 16, 29, 43, 53], but may have limited expressiveness and high memory and computation requirements, respectively. Moreover, fully implicit representations-based approaches  [8, 35] have also been proposed, but their slow querying and sampling make them difficult to use in training. Several works have also proposed that utilize a hybrid approach  [45, 7, 14, 46].
@@ 3 Method
In this work, we aim to leverage a 3D GAN to enable novel attribute editing driven by natural language prompts that are 3D-aware and view-consistent. We use frozen 3D-aware StyleGAN called GMPI [50] to edit 3D images and guide its latent space ($\mathcal{W}$) via. our proposed latent attribute editor (LAE) in the target attribute direction by utilizing CLIP [32]). Our method is significantly efficient as it only requires training of attribute-specific style tokens and linear-layer-based style mappers, which can be plugged in with different 3D generative models.
@@ 3.1 Problem Formulation
Given an input latent code $z$, an attribute editing instruction $A^{i}$ and a target camera pose $p_{t}$; our goal is to generate multiplane representations $\mathcal{M}$ that can be used to render 3D-aware and view-consistent images having the given attributes. A multiplane image can be represented as $(C_{i},\alpha_{i},d_{i})$ for $L$ fronto-parallel planes, where $C\in\mathbb{R}^{H\times H\times 3}$ denotes color texture, $\alpha_{i}\in[0,1]^{H\times H\times 1}$ and $d_{i}\in\mathbb{R}$ denotes alpha maps and depth for corresponding plane (distance from a camera). We use GMPI [50], which simplifies this task to the generation of a single color texture image across all planes along with alpha maps. The alpha maps along with color texture $C_{i}$ are then fed to a renderer $R$ to generate an image at a specific camera angle. However, we also want to edit these images to have an arbitrary attribute $A^{i}$ specified by natural language text. Hence, the goal of our generator $f_{G}$ is to generate an RGB image $C$ with a particular attribute and corresponding alpha maps given a latent code $z$, a textual instruction $A^{i}$ and depth of the planes $d_{i}$: $M=\{C,\{\alpha_{i},...,\alpha_{L}\}\}=f_{G}(z,A^{i},\{d_{1},...,d_{L}\})$.
@@ 3.2 Overview of Proposed Pipeline
Recent 3D face editing methods mostly suffer from several limitations, such as rigidity and time consumption, primarily because they rely on predefined attribute classes. This dependence requires training pre-trained attribute classifiers on large datasets, which can be computationally intensive. As a result, these methods struggle to adapt to novel attributes in real-time scenarios, limiting their practical utility for 3D-aware editing. To address these challenges, we introduce a learnable prompt-based attribute editor $LAE$ module within the GMPI framework and 3D-aware attribute editing, identity and pose preservation losses for synthesizing and editing face images with prompt controllable attributes (e.g., hair color, style, expressions) while maintaining the view-consistency across various target poses. Further, the proposed $LAE$ module can also be integrated into other state-of-the-art 3D generation methods, enabling the editing capabilities while maintaining the identity and the view consistency across multiple camera poses.
@@
Our proposed framework is trained end-to-end using the proposed 3D-aware attribute editing loss, prompt-based editing loss, and 3D-aware identity and pose preservation losses. While prompt-based editing loss ($\mathcal{L}_{dclip}$, $\mathcal{L}_{sc}$) enables controllable editing of attributes in the generated face images, the identity preservation loss ($\mathcal{L}_{id}$, $\mathcal{L}_{idvc}$, $\mathcal{L}_{latent}$, and $\mathcal{L}_{\alpha}$) strives to maintain the identity and camera pose. In the following, we explain our language-driven attribute editor (LAE) and losses designed to preserve the identity and 3D consistency of the generated images.
@@ 3.3 Latent Attribute Editor (LAE) with Text Driven Editing
While the latent space of StyleGAN [21] has shown to be fairly disentangled, it still requires finding editing direction corresponding to an attribute. Since our goal is text-driven editing, an important question is: how can we effectively learn and extract information from the attribute prompt $A_{i}$ so that the network can generate images with the desired attributes? Inspired by [30], we propose a text-driven latent attribute editor (LAE) consisting of style tokens and style mappers trained on Directional Clip and style token contrastive losses.
@@
While $\Delta v$ represents style, it can not be directly fed to StyleGAN as it is incompatible with its latent space. To transform $\Delta v$ to $\Delta w$, we use linear mappers  [21]: $\Delta w=M(\Delta v)$. Given a random latent vector $z$ as an input, a pretrained $f_{Map}$ aims to get a latent code $w$. These latent codes $w$ are split into three groups (coarse $w_{c}$, middle $w_{m}$ and fine $w_{f}$) in the StyleCLIP [30]. The Style Mapper consists of three sub-networks for coarse, middle, and fine features, each consisting of single linear layers. The style mapper takes the output of the CLIP text encoder ($\Delta v$) and latent code $w$ and outputs editing direction $\Delta w$ for StyleGAN. The three levels of Style Mapper $M$ can be formulated as:
@@
$\hat{w}^{i}=w^{i}+\mathbf{M}(w^{i},\Delta v^{i})$. The $\hat{w}$ are passed as a input to the $G(.)$, which generates multiplane image (MPI) representation $D=\{C,\{\alpha_{1}\cdots\alpha_{L}\}\}$, these MPIs $D$ along with the target camera pose $p_{t}$ are feed in the MPI Renderer $R$ to get the final generated image: ${I_{p_{t}}^{i}}=R(G_{t}(\hat{{w}}^{i},\theta),p_{t})$
We train style tokens and style mappers on a few hundred synthetic images. Our method is data-free and extremely efficient to train. Since it is language-driven and efficient, it can be used on the fly to add and edit arbitrary new attributes. Moreover, a single set of mappers is sufficient for editing a large number of attributes by only adding additional style tokens.
Figure 3: Our method’s performance in face editing is compared qualitatively with both (GMPI [50]) and the state-of-the-art PREIM3D [24] across various camera angles and attributes. The following attributes were used for comparison: young for age, blond for hair color, and happy for emotion. Additionally, to showcase our method’s ability to enable the editing of novel attributes, the final column presents results obtained from custom prompts. Our method accurately maintains camera poses in relation to GMPI and demonstrates superior identity preservation and editing capability compared to PREIM3D.
Directional CLIP Loss for Attribute Editing. One simple method for guiding a generated image $I_{p_{t}}^{i}$ with the textual prompt $P_{A}^{i}$ is to align it with a target text prompt’s semantic is to use a CLIP-based image manipulation approach [30]. This approach involves minimizing a global clip loss function that is formulated as a way to achieve this alignment. However, such a global clip loss led to low diversity and corrupt outputs. To address these issues, we utilize enhanced Directional CLIP loss $L_{dclip}$ [48], which differs from the approach used in Gal et al. [12] as it uses attribute-specific prompts and generates multi-view images on-the-fly, rather than relying on fixed, manually designed prompts and single-view images. For a given latent code $w^{i}$, we compute the direction of the source and target image pair at two different camera poses $p_{t_{1}}$ and $p_{t_{2}}$,
@@
Contrastive Learning for Simultaneous Learning of Multiple Style Tokens Our approach employs $n$ distinct style tokens, each dedicated to representing a unique attribute. Since these tokens share a style mapper, it is possible for the style tokens to converge to a common orthogonal point in the text embedding space, which is common to all the facial attributes. To prevent this, we introduce a novel style token contrastive loss or $\mathcal{L}_{sc}$. This loss works by minimizing the similarity learned style tokens. Specifically, for a set of $n$ attribute-specific style tokens $\{P^{1}_{A},P^{2}_{A}\cdots P^{n}_{A}\}$, the style token contrastive loss is defined as,
@@
Given the latent code $\hat{w}$, our goal is to generate 3D-aware images with the different attribute prompt $P_{A}$ from a single model. To address this objective, we examine the following questions. First, How to preserve the identity in the image across different camera poses $p_{t}$ and different attributes. Second, How can we prevent excessive deviation of $\alpha-$maps during training, particularly for attributes like expression and hairstyle?
The first issue is addressed by enforcing consistency in the multi-view output across different attributes and camera poses. Finally, $\alpha-$maps preserving regularization is introduced to avoid the drastic change in the alpha maps during attribute editing. To preserve identity and pose, we utilize a few 3D-aware losses. The textual loss ensures the final generated image $I_{p_{t}}$ aligns with the text prompts and is realistic, i.e., artifact-free. However, it does not constrain the identity with respect to the original image generated from $G_{o}$ and the generated image with the different camera poses and different attributes.
Table 2: We compare our method with the 3D-Inv [27], PixelNeRF [6] and PRIEM3D [24] for attribute altering (AA) and attribute dependency (AD) metrics across various attributes following PREIM3D’s protocol. Our method significantly outperforms in both measures. Here, "NA" denotes that the model doesn’t have a trained classifier to edit at the particular attribute, whereas our model can generate at any novel attribute.
@@
Identity Preservation. Identity loss is utilized to ensure identity consistency across different camera poses and before and after the attribute change. The identity loss aims to preserve the identity of original and modified images at a fixed front camera pose $p_{o}$. It is defined as follows:
@@
where the first term is attribute modified image at the camera, pose $p_{t}$ and the second term is the unmodified image at the same camera pose, $\cos(\cdot)$ is the cosine similarity, and $AF(\cdot)$ is a pretrained ArcFAce Network [10] for face recognition. Here, we set the $p_{t}$ as the frontal angle.
@@
$\mathcal{L}_{idvc}=\sum_{i=1}^{K}\sum_{j=1,{i\neq j}}^{K-1}1-\cos(AF(R(G_{t}(\hat{w}^{i}),p_{t_{1}}),R(f_{G_{t}}(\hat{w}^{j}),p_{t_{2}}))),$
where $p_{t_{1}}$ and $p_{t_{2}}$ are two different camera poses sampled randomly. $\mathcal{L}_{idvc}$ ensures identity consistency between different camera poses and different text prompts.
Figure 4: Our method edits various attributes while preserving both identity and camera poses.
@@
Table 3: Comparison of depth and pose accuracy between our and state-of-the-art 3D generation methods. Our method demonstrates depth and pose accuracy close to the baseline state-of-the-art 3D generation method.

### v2

@@ Abstract
Drawing upon StyleGAN’s expressivity and disentangled latent space, existing 2D approaches employ textual prompting to edit facial images with different attributes. In contrast, 3D-aware approaches that generate faces at different target poses require attribute-specific classifiers, learning separate model weights for each attribute, and are not scalable for novel attributes. In this work, we propose an efficient, plug-and-play, 3D-aware face editing framework based on attribute-specific prompt learning, enabling the generation of facial images with controllable attributes across various target poses. To this end, we introduce a text-driven learnable style token-based latent attribute editor (LAE). The LAE harnesses a pre-trained vision-language model to find text-guided attribute-specific editing direction in the latent space of any pre-trained 3D-aware GAN. It utilizes learnable style tokens and style mappers to learn and transform this editing direction to 3D latent space. To train LAE with multiple attributes, we use directional contrastive loss and style token loss. Furthermore, to ensure view consistency and identity preservation across different poses and attributes, we employ several 3D-aware identity and pose preservation losses. Our experiments show that our proposed framework generates high-quality images with 3D awareness and view consistency while maintaining attribute-specific features. We demonstrate the effectiveness of our method on different facial attributes, including hair color and style, expression, and others.
@@ 1 Introduction
StyleGAN [21, 22] has demonstrated exceptional capabilities in generating unconditional, photorealistic 2D images. The disentangled properties of the learned latent space of StyleGAN have facilitated attribute-specific realistic image editing by finding and modifying semantic editing directions  [1, 37, 15, 42].
@@
Furthermore, recent studies have harnessed foundational vision-language models (e.g., CLIP  [33]) to facilitate text-guided attribute editing [31].
@@
Recently, StyleGAN-based architecture has also been adapted for 3D-aware and view-consistent image generation [51, 7, 14, 28, 40]. For instance, GMPI [51] has modified 2D StyleGAN by introducing an alpha branch to learn alpha maps to generate 3D-aware multiplane images efficiently. StyleNeRF [14] integrates the neural radiance field (NeRF) into a style-based generator and EG3D [7] which incorporate triplane representation along with style mapping network to generate view consistent images. However, compared to 2D images, achieving 3D-aware image editing encounters several challenges, as it requires not only maintaining consistency between the edited and original images but also ensuring view consistency across different poses while preserving the facial identity.
Building on 3D GANs, some works have introduced attribute editing methods that can manipulate a set of attributes in 3D [28, 40, 35, 25]. However, these methods are limited to editing a predetermined set of attributes as they require training attribute classifiers for every new attribute. This training takes several hours on large datasets of training for every attribute. Moreover, these methods struggle with maintaining identity and view consistency across a wide variety of camera angles. Finally, due to the use of pre-trained classifiers, novel attribute editing is expensive and limited. A more natural way is to use textual prompts that not only significantly extend the range of manipulations but also provide a more intuitive form of interaction that closely resembles humans [31].
@@
In this work, we introduce a plug-and-play module designed to efficiently integrate text-guided editing of novel attributes into 3D-aware GANs while preserving both the 3D pose and identity. Our proposed module, Learned Attribute Editor (LAE), finds semantic editing directions specified by a natural language prompt, thereby facilitating text-guided manipulation of novel attributes. The language prompt specifying the target attribute is concatenated with learnable style tokens and passed through a frozen CLIP text encoder. These prompt encodings are then passed through a set of style mappers that learn to transform text-specified modifications to the latent space of 3D-aware GAN. The combination of learnable style tokens and vision-language models makes the editing process extremely efficient (from hours to a few minutes), as LAE does not need to train classifiers on large datasets for every new attribute.
@@
Training the LAE module with language supervision poses challenges, as standard CLIP loss [33] may struggle to converge in the 3D latent space. To overcome this issue, we introduce a 3D-aware DCLIP loss [12] which utilizes multi-view images generated on-the-fly. To improve the view consistency and preserve multi-view identity, we introduce novel losses, including style token contrastive loss, IDVC, and several 3D-aware pose and identity preservation losses. These losses enable maintaining the distinctive characteristics of each attribute, aligning the CLIP-space direction with learnable text features and the multi-view generated images, and ensuring consistency across different camera poses. We primarily employ GMPI [51], which adapts StyleGAN to generate color images and corresponding alpha maps. We also demonstrate our method’s plug-and-play nature by integrating it with several different 3D-aware GANs. Our experimental results validate the efficacy of our proposed method, showcasing its superior generative performance in both qualitative and quantitative ways. We present a representative sample of editing made by our method at random camera poses in Figure 1.
@@ 2 Related Works
Generative 3D-aware image synthesis and manipulation. Extending the capabilities of 2D GANs to 3D settings has gained significant attention recently. These methods rely on a combination of 3D-structure aware inductive bias in the generator and utilization of neural rendering engines to get view-consistent results. These approaches include Mesh-based appraoches [27, 41], Voxel-based GANs [11, 16, 30, 44, 54], but may have limited expressiveness and high memory and computation requirements, respectively. Moreover, fully implicit representations-based approaches  [8, 36] have also been proposed, but their slow querying and sampling make them difficult to use in training. Several works have also proposed that utilize a hybrid approach  [46, 7, 14, 47, 24].
@@ 3 Method
In this work, our goal is to leverage a 3D GAN to enable novel attribute editing driven by natural language prompts that are 3D-aware and view-consistent. We use frozen 3D-aware StyleGAN called GMPI [51] to edit 3D images and guide its latent space ($\mathcal{W}$) via. our proposed latent attribute editor (LAE) in the target attribute direction by utilizing CLIP [33]). Our method is significantly efficient as it only requires training of attribute-specific style tokens and linear-layer-based style mappers, which can be plugged in with different 3D generative models.
@@ 3.1 Problem Formulation
Given an input latent code $z$, an attribute editing instruction $A^{i}$ and a target camera pose $p_{t}$, our goal is to generate multiplane representations $\mathcal{M}$ that can be used to render 3D-aware and view-consistent images having the given attributes. A multiplane image can be represented as $(C_{i},\alpha_{i},d_{i})$ for $L$ fronto-parallel planes, where $C\in\mathbb{R}^{H\times H\times 3}$ denotes color texture, $\alpha_{i}\in[0,1]^{H\times H\times 1}$ and $d_{i}\in\mathbb{R}$ denotes alpha maps and depth for corresponding plane (distance from a camera). We use GMPI [51], which simplifies this task to the generation of a single color texture image across all planes along with alpha maps. The alpha maps along with color texture $C_{i}$ are then fed to a renderer $R$ to generate an image at a specific camera angle. However, we also want to edit these images to have an arbitrary attribute $A^{i}$ specified by natural language text. Hence, the goal of our generator $f_{G}$ is to generate an RGB image $C$ with a particular attribute and corresponding alpha maps given a latent code $z$, a textual instruction $A^{i}$ and depth of the planes $d_{i}$: $M=\{C,\{\alpha_{i},...,\alpha_{L}\}\}=f_{G}(z,A^{i},\{d_{1},...,d_{L}\})$.
@@ 3.2 Overview of Proposed Pipeline
Recent 3D face editing methods mostly suffer from limitations such as rigidity and time consumption, primarily because they rely on predefined attribute classes. This dependence requires training pre-trained attribute classifiers on large datasets, which can be computationally intensive. As a result, these methods struggle to adapt to novel attributes in real-time scenarios, limiting their practical utility for 3D-aware editing. To address these challenges, we introduce a learnable prompt-based attribute editor $LAE$ module within the GMPI framework and 3D-aware attribute editing, identity, and pose preservation losses for synthesizing and editing face images with prompt controllable attributes (e.g., hair color, style, expressions) while maintaining the view-consistency across various target poses. Further, the proposed $LAE$ module can also be integrated into other state-of-art 3D generation methods, enabling the editing capabilities while maintaining the identity and the view consistency across multiple camera poses.
@@
Our proposed framework is trained end-to-end using the proposed 3D-aware attribute editing loss, prompt-based editing loss, and 3D-aware identity and pose preservation losses. While prompt-based editing loss ($\mathcal{L}_{dclip}$, $\mathcal{L}_{sc}$) enables controllable editing of attributes in the generated face images, the identity preservation loss ($\mathcal{L}_{id}$, $\mathcal{L}_{idvc}$, $\mathcal{L}_{latent}$, and $\mathcal{L}_{\alpha}$) strives to maintain the identity and camera pose. In the following, we first explain our language-driven attribute editor (LAE) and then losses that are designed to preserve the identity and 3D consistency of the generated images.
@@ 3.3 Latent Attribute Editor (LAE) with Text Driven Editing
While the latent space of StyleGAN [21] has shown to be fairly disentangled, it still requires finding editing direction corresponding to an attribute. Since our goal is text-driven editing, an important question is: how can we effectively learn and extract information from the attribute prompt $A_{i}$ so that the network can generate images with the desired attributes? Inspired by [31], we propose a text-driven latent attribute editor (LAE), which consists of style tokens and style mappers and is trained on Directional Clip and style token contrastive losses.
@@
While $\Delta v$ represents style, it can not be directly fed to StyleGAN as it is not compatible with its latent space. To transform $\Delta v$ to $\Delta w$, we use linear mappers  [21]: $\Delta w=M(\Delta v)$. Given a random latent vector $z$ as an input, a pre-trained mapping network $f_{Map}$ aims to get a latent code $w$. These latent codes $w$ are split into three groups (coarse $w_{c}$, middle $w_{m}$ and fine $w_{f}$) in the StyleCLIP [31]. The Style Mapper consists of three sub-networks for coarse, middle, and fine features, and each one consists of single linear layers. The style mapper takes the output of the CLIP text encoder ($\Delta v$) and latent code $w$ and outputs editing direction $\Delta w$ for StyleGAN. The three levels of Style Mapper $M$ can be formulated as:
@@
$\hat{w}^{i}=w^{i}+\mathbf{M}(w^{i},\Delta v^{i})$. The $\hat{w}$ are passed as a input to the $G(.)$, which generates multiplane image (MPI) representation $D=\{C,\{\alpha_{1}\cdots\alpha_{L}\}\}$, these MPIs $D$ along with the target camera pose $p_{t}$ are feed in the MPI Renderer $R$ to get the final generated image through generator $G_{t}$: ${I_{p_{t}}^{i}}=R(G_{t}(\hat{{w}}^{i},\theta),p_{t})$
Our method is data-free and extremely efficient to train. Since it is language-driven and efficient, it can be used on the fly to add and edit arbitrary new attributes. Moreover, a single set of mappers is sufficient for editing a large number of attributes by only adding additional style tokens.
Figure 3: Our method’s performance in face editing is compared qualitatively with both (GMPI [51]) and state-of-the-art PREIM3D [25] across various camera angles and attributes. The following attributes were used for comparison: young for age, blond for hair color, and happy for emotion. Additionally, to showcase our method’s ability to enable the editing of novel attributes, the final column presents results obtained from custom prompts. Our method not only accurately maintains camera poses about GMPI but also demonstrates superior identity preservation and editing capability in comparison to PREIM3D.
Directional CLIP Loss for Attribute Editing. A simple method for guiding a generated image $I_{p_{t}}^{i}$ with the textual prompt $P_{A}^{i}$ is to align it with a target text prompt’s semantic to use a CLIP-based image manipulation approach [31]. This approach involves minimizing a global clip loss function that is formulated as a way to achieve this alignment. However, such a global clip loss led to low diversity and corrupt outputs. To address these issues, we utilize enhanced Directional CLIP loss $L_{dclip}$ [49], which differs from the approach used in Gal et al. [12] as it uses attribute-specific prompts and generates multi-view images on-the-fly, rather than relying on fixed, manually designed prompts and single-view images. For a given latent code $w^{i}$, we compute the direction of the source and target image pair at two different camera poses $p_{t_{1}}$ and $p_{t_{2}}$,
@@
Contrastive Learning for Simultaneous Learning of Multiple Style Tokens Our approach employs $n$ distinct style tokens, each dedicated to representing a unique attribute. Since these tokens share a style mapper, the style tokens can converge to a common orthogonal point in the text embedding space, which is common to all facial attributes. To prevent this, we introduce a novel style token contrastive loss or $\mathcal{L}_{sc}$. This loss works by minimizing the similarity learned style tokens. Specifically, for a set of $n$ attribute-specific style tokens $\{P^{1}_{A},P^{2}_{A}\cdots P^{n}_{A}\}$, the style token contrastive loss is defined as,
@@
Given the latent code $\hat{w}$, our goal is to generate 3D-aware images with the different attribute prompt $P_{A}$ from a single model. To address this objective, we examine the following questions. First, How to preserve the identity of the image across different camera poses $p_{t}$ and different attributes. Second, How to prevent excessive deviation of $\alpha-$maps during training, particularly for attributes like expression and hairstyle?
The first issue is addressed by enforcing consistency in the multi-view output across different attributes and camera poses. Finally, $\alpha-$maps preserving regularization is introduced to avoid the drastic change in the alpha maps during attribute editing. To preserve identity and pose, we utilize a few 3D-aware losses. The textual loss ensures the final generated image $I_{p_{t}}$ aligns with the text prompts and is realistic, i.e., artifact-free. However, it does not constrain the identity of the image concerning the original image generated from $G_{o}$ and the generated image with the different camera poses and different attributes.
Table 2: We compare our method with the 3D-Inv [28], PixelNeRF [6] and PRIEM3D [25] for attribute altering (AA) and attribute dependency (AD) metrics across various attributes following PREIM3D’s protocol. Our method significantly outperforms in both measures. Here "NA" denotes that the model doesn’t have a trained classifier to edit the particular attribute, whereas our model can generate any novel attribute.
@@
Identity Preservation. To ensure identity consistency across different camera poses and before and after the attribute change, identity loss is utilized. The identity loss aims to preserve the identity of original and modified images at a fixed front camera pose $p_{o}$. It is defined as follows:
@@
where the first term is attribute modified image at the camera, pose $p_{t}$ and the second term is the unmodified image at the same camera pose, $\cos(\cdot)$ is the cosine similarity, and $AF(\cdot)$ is a pre-trained ArcFAce Network [10] for face recognition. Here, we set the $p_{t}$ as the frontal angle.
@@
$\mathcal{L}_{idvc}=\sum_{i=1}^{K}\sum_{j=1,{i\neq j}}^{K-1}1-\cos(AF(R(G_{t}(\hat{w}^{i}),p_{t_{1}}),R(G_{t}(\hat{w}^{j}),p_{t_{2}}))),$
where $p_{t_{1}}$ and $p_{t_{2}}$ are two different camera poses sampled randomly. $\mathcal{L}_{idvc}$ ensures the identity consistency between different camera poses and different text prompts.
@@
Table 3: Comparison of depth and pose accuracy between our method and the state-of-the-art 3D generation method. Our method demonstrates depth and pose accuracy close to that of the baseline state-of-the-art 3D generation method.